# Trojan Cleansing with Neural Collapse

### v2

@@ 4.3.3 Adaptive Attack
Consider a scenario where the adversary is aware that trojan attacks undermine the phenomenon of Neural Collapse. Thus, they train the model with the final layer’s weights fixed as an ETF, effectively replicating the conditions under which Neural Collapse occurs perfectly. ETF-FT demonstrates robust performance even in the face of adaptive attacks. As evidenced in Tab. 6, the ASR is only 90.65, which is lower than the ASR observed in other similar settings. This supports our observation that trojan attacks are incompatible with the Neural Collapse mechanism. Notably, ETF-FT not only removes the trojans but also enhances the model’s Accuracy.
@@
Table 6: Adaptive Attack: Performance of ETF-FT on trojaned model trained with final layer fixed as ETF. Results are reported on 10% Badnets trojaned CIFAR-10 using PreAct-ResNet18, averaged over 10 random trials (mean $\pm$ std).
@@ 4.3.4 Cleansing Without Overtraining
Neural Collapse occurs most strongly in models trained well past zero training error. In line with this, our other experiments used models trained for 350 epochs, despite reaching zero training error at approximately 120 epochs. However, in practical scenarios, models may not always be overtrained. To examine ETF-FT’s reliance on model overtraining we also test it on models trained only until the training error reaches zero. As shown in Tab. 7, ETF-FT maintains strong performance in this case, achieving comparable results to its performance on overtrained models.
@@
Table 7: No Overtraining: Performance of ETF-FT on trojaned model only trained for 120 epochs. Results are reported on 10% Badnets trojaned CIFAR-10 using PreAct-ResNet18, averaged over 10 random trials (mean $\pm$ std).
@@ Appendix A More Cleansing Results
NCleanse[29] $\textbf{88.43}\pm 0.29$ $\textbf{1.19}\pm 0.27$
NAD[17] $83.41\pm 1.87$ $3.17\pm 0.72$
NPD[37] $83.10\pm 3.64$ $5.97\pm 4.59$
FT-SAM[36] $87.20\pm 0.54$ $7.17\pm 3.69$
I-BAU[33] $85.97\pm 1.43$ $49.55\pm 12.15$
@@
Table 8: Data Imbalance: Comparison of cleansing algorithms with 1% imbalanced clean training data on BadNet trojaned ResNet models(%). Results are averaged over 5 random trials with mean±$\pm$±std.

### v3

@@ 4.3.3 Adaptive Attack
Consider a scenario where the adversary is aware that trojan attacks undermine the phenomenon of Neural Collapse. Thus, they train the model with the final layer’s weights fixed as an ETF, effectively replicating the conditions under which Neural Collapse occurs perfectly. ETF-FT demonstrates robust performance even in the face of adaptive attacks. As evidenced in section 4.3.3, the ASR is only 90.65, which is lower than the ASR observed in other similar settings. This supports our observation that trojan attacks are incompatible with the Neural Collapse mechanism. Notably, ETF-FT not only removes the trojans but also enhances the model’s Accuracy.
@@
[…]
tableAdaptive Attack: Performance of ETF-FT on models trained with final layer fixed as ETF. Results are on BadNets-trojaned CIFAR-10 using PreAct-ResNet18, averaged over 10 trials.
@@ 4.3.4 Cleansing Without Overtraining
Neural Collapse occurs most strongly in models trained well past zero training error. In line with this, our other experiments used models trained for 350 epochs, despite reaching zero training error at approximately 120 epochs. However, in practical scenarios, models may not always be overtrained. To examine ETF-FT’s reliance on model overtraining we also test it on models trained only until the training error reaches zero. As shown in section 4.3.4, ETF-FT maintains strong performance in this case, achieving comparable results to its performance on overtrained models.
@@
[…]
tableNo Overtraining: Performance of ETF-FT on trojaned model only trained for 120 epochs. Results are on BadNets-trojaned CIFAR-10 using PreAct-ResNet18, averaged over 10 trials.
4.3.5 Complex Datasets
We further assess the robustness of ETF-FT on two challenging benchmarks—CIFAR-100 (100 classes) and GTSRB (43 traffic-sign classes with non-uniform image sizes)—to probe generalization under greater class granularity and real-world variability. As summarized in Table 7, ETF-FT consistently maintains a high clean accuracy while substantially reducing ASR in both datasets, maintaining a high competency among the cleansing methods.
Method CIFAR-100 GTSRB
ACC ASR ACC ASR
Original 70.69 87.25 95.96 89.17
FT 70.80 $\pm$ 0.14 47.02 $\pm$ 19.59 97.39 $\pm$ 0.16 0.01 $\pm$ 0.02
NCleanse [neuralCleanse] 69.96 $\pm$ 0.54 0.05 $\pm$ 0.03 97.05 $\pm$ 0.27 0.00 $\pm$ 0.01
NAD[NeuralAttentionDistillation] 70.61 $\pm$ 0.21 42.18 $\pm$ 25.56 97.10 $\pm$ 0.27 0.00 $\pm$ 0.01
NPD[zhu2024neural] 60.39 $\pm$ 0.72 0.21 $\pm$ 0.11 95.29 $\pm$ 0.30 0.58 $\pm$ 0.34
FT-SAM[zhu2023enhancing] 68.51 $\pm$ 0.30 16.33 $\pm$ 15.13 97.98 $\pm$ 0.15 0.01 $\pm$ 0.01
I-BAU[zeng2021adversarial] 68.63 $\pm$ 0.74 3.34 $\pm$ 3.81 95.06 $\pm$ 0.67 0.05 $\pm$ 0.06
ETF-FT (ours) 66.09 $\pm$ 0.39 0.36 $\pm$ 0.13 97.90 $\pm$ 0.19 0.00 $\pm$ 0.00
Table 7: CIFAR100 & GTSRB: Comparison of cleansing methods on on BadNet attacked PreAct-ResNet18 models with 5% clean data. Averaged over 10 trials.
@@ Appendix A More Cleansing Results
NCleanse[neuralCleanse] $\textbf{88.43}\pm 0.29$ $\textbf{1.19}\pm 0.27$
NAD[NeuralAttentionDistillation] $83.41\pm 1.87$ $3.17\pm 0.72$
NPD[zhu2024neural] $83.10\pm 3.64$ $5.97\pm 4.59$
FT-SAM[zhu2023enhancing] $87.20\pm 0.54$ $7.17\pm 3.69$
I-BAU[zeng2021adversarial] $85.97\pm 1.43$ $49.55\pm 12.15$
@@
[…]
Appendix B Standard Deviation of Table 4
We append the standard deviation of table 4 here for reference.
Method WaNet BPP Input-aware
ACC ASR ACC ASR ACC ASR
Original
FT 0.09 0.21 0.23 0.35 0.28 0.28
NCleanse[neuralCleanse] 0.76 47.30 3.67 39.21 2.82 28.86
NAD[NeuralAttentionDistillation] 0.31 0.23 2.57 5.04 4.39 3.74
NPD[zhu2024neural] 4.22 2.40 5.74 11.56 1.88 2.02
FT-SAM[zhu2023enhancing] 0.13 0.11 0.65 0.50 0.49 1.48
I-BAU[zeng2021adversarial] 0.38 3.51 0.73 31.21 0.57 11.70
ETF-FT (ours) 0.13 0.19 0.15 9.54 0.18 0.26
Table 9: Standard deviation for the comparison of cleansing methods on WaNet, BPP, and Input-aware attacked CIFAR-10 models. Results are calculated over 10 trials.
Appendix C Code and Instructions
Our implementation is integrated into the BackdoorBench [wu2022backdoorbench] repository. The core implementation of our method is provided in the file etf-ft.py, which should be placed under BackdoorBench/defense/. The directory etf-ft/ contains configuration files and hyperparameter settings used for our experiments and should be placed under BackdoorBench/config/defense/. Note that ETF-FT relies on the AdamW optimizer and ExponentialLR scheduler, which must be added to the BackdoorBench environment if not already available. Once configured, ETF-FT can be executed using the same command interface as other defense methods in the framework.
For experiments involving the Instagram filter trigger, the transformation is implemented in filter.py, based on the module proposed in [trojai_instagram_xforms], and should be placed in BackdoorBench/utils/bd_img_transform/.
Appendix D Experimental Details
ETF-FT employs the AdamW optimizer in conjunction with an ExponentialLR learning rate scheduler. The specific learning rate and decay factor (gamma) are tuned based on the model architecture.
For PreAct-ResNet18 (on CIFAR-10/CIFAR-100/GTSRB), we use:
  • Learning rate: 0.0001
  • Learning rate scheduler: ExponentialLR
  • gamma: 0.95
  • Optimizer: AdamW
For ViT on CIFAR-10, we use:
  • Learning rate: 0.001
  • Learning rate scheduler: ExponentialLR
  • gamma: 0.85
  • Optimizer: AdamW